# Graphon Mixtures

### v1

@@ B.1 A $(U,W)$ mixture example
When $\frac{n_{s_{i}}}{n_{d_{i}}}\to c\in\mathbb{R}$ the graph sequence $\{G_{n_{i}}\}_{i}$ is dense (Lemma 3.4), i.e., the edge density of $G_{n_{i}}$ does not go to zero. Figures 9 and 9 show degree distributions of two examples when $\frac{n_{s_{i}}}{n_{d_{i}}}=1$ and $\frac{n_{s_{i}}}{n_{d_{i}}}=3$. When the ratio $\frac{n_{s_{i}}}{n_{d_{i}}}$ converges to $c\leq 1$, depending on $W$ and $U$, the largest degree may be contributed by the dense part. In Figure 9 the largest degree is actually produced by $U$, however, it is so close to the degrees generated by $W$ that we cannot distinguish the effect of $U$ in this example.
When the ratio $\frac{n_{s_{i}}}{n_{d_{i}}}$ converges to $c>1$, for large enough $i$ the sparse part contributes to highest-degree nodes, which are anomalous. In Figure 9 two anomalous nodes are contributed by $U$ for each pair of $n_{s_{i}}$ and $n_{d_{i}}$.
Figure 6: When $\frac{n_{s_{i}}}{n_{d_{i}}}=1$
@@
If $\{G_{n_{i}}\}_{i}$ is sparse, then $n_{s_{i}}$ grows faster than $n_{d_{i}}$ making $\frac{n_{s_{i}}}{n_{d_{i}}}$ go to infinity (Lemma 3.4). We used $n_{s_{i}}=\lceil\sqrt{5i}\times n_{d_{i}}\rceil$ for this experiment and Figure 9 shows the degree distributions. Two anomalous nodes contributed by $U$ can be seen in 3 of the 4 subplots.
@@ B.2 Expectations
B.3 Edge density of $(U,W)$-mixture graphs
@@
Let $\{G_{n_{i}}\}_{i}$ be a sequence of sparse $(U,W)$-mixture graphs (Definition 3.3) with dense and sparse parts of $G_{n_{i}}$ being $G_{d_{i}}$ and $G_{s_{i}}$ respectively. Let $\bm{p}=(p_{1},p_{2},\ldots)$ be the mass-partition (Definition 3.2) associated with $U$ and suppose $\bm{p}$ has $k$ non-zero elements. Recall $G_{s_{i}}$ is a union of disjoint stars and isolated edges. Let $\tilde{q}_{j,i}$ be the degree of the star in $G_{s_{i}}$ corresponding to $p_{j}\neq 0$. Then, if we consider $G_{n_{i}}$ to be the disjoint union of $G_{d_{i}}$ and $G_{s_{i}}$ the top $k$ degrees denoted by deg $\text{deg }v_{(1\ldots k)}$ satisfies
@@
giving the result. ∎
B.4 When $\sum_{i}p_{i}<1$
@@ D.2 Estimating $k$
Let $\{G_{n_{i}}\}_{i}$ be a sequence of sparse $(U,W)$-mixture graphs (Definition 3.3) with dense and sparse parts $G_{d_{i}}$ and $G_{s_{i}}$ respectively. Let $\bm{p}=(p_{1},p_{2},\ldots)$ be the mass-partition (Definition 3.2) associated with $U$ with $k$ finite partitions.Let $\tilde{q}_{j,i}$ be the degree of the star in $G_{s_{i}}$ corresponding to $p_{j}\neq 0$. Let $q_{j,i}$ denote the degree of the corresponding vertex in $G_{n_{i}}$. Let deg $\text{deg }v_{(\ell)}$ denote the $\ell$th largest degree in $G_{n_{i}}$. Suppose deg $\text{deg }{v}_{(\ell)}=q_{\ell,i}$ for $\ell\in\{1,\ldots,k\}$ and let $\alpha\geq 0$ be a constant. Then for any $\ell\in\{1,\ldots,(k-1)\}$
@@
[…]
@@
[…]
@@
[…]
@@
where we have used equation (113). In the logscale the difference between $k$th highest vertex and $(k+1)$st highest vertex is larger than the $\ell$th highest vertex and the $(\ell+1)$st highest vertex for $\ell<k$. That is, in the log scale, successive difference of the top $k$ degrees is highest at $k$. From Lemma D.4 we know that for degrees in the dense part $G_{d_{i}}$, the ratios of successive degrees converge to 1, when we exclude small degrees. As we are taking ordered values $\text{deg }_{G_{n_{i}}}v_{(\ell)}/\text{deg }_{G_{n_{i}}}v_{(\ell+1)})\approx p_{\ell}/p_{\ell+1}>1$. The ratio $\text{deg }_{G_{n_{i}}}v_{(k)}/\text{deg }_{G_{n_{i}}}v_{(k+1)}\approx m_{s_{i}}p_{k}/n_{d_{i}}D_{\max}$ where $D_{\max}=\max D(x)$. Thus, the log difference $\log\text{deg }_{G_{n_{i}}}v_{(k)}-\log\text{deg }_{G_{n_{i}}}v_{(k+1)}$ is much greater than 1 for $i>I_{0}$. Therefore we have
@@ Appendix E Estimating $U$ for infinite partitions
We turn out attention to the term $\hat{k}_{i}$ We note that $\hat{k}_{i}\in O(m_{s_{i}}/n_{d_{i}})$. The sequence $\{\hat{k}_{i}\}_{i}$ increases at a slower rate compared to $m_{s_{i}}/n_{d_{i}}$ depending on the rate $\{p_{j}\}_{j}$ decreases. For $\sum_{j}p_{j}$ to add to 1, $p_{j}$ cannot decrease like $c/j$ for come constant $c$, because $\sum_{j}c/j$ diverges to infinity. Hence $p_{j}\approx c/j^{1+\alpha}$ is an option. The estimate $\hat{k}_{i}$ satisfies
@@
$\lim_{i\to\infty}\frac{1}{m_{s_{i}}}\text{Var}(Y_{i})=\sum_{\ell=1}^{\hat{k}_{i}}p_{\ell}\left(1-\sum_{\ell=1}^{\hat{k}_{i}}p_{\ell}\right)\,.\$
@@
[…]
@@ F.1.1 Degree prediction
As a baseline comparison method we used a scale-free graph property discussed in Bollobás et al. (2001). For a scale-free graph with $n$ nodes the maximum degree behaves like $\Theta(\sqrt{n})$. Suppose $\text{deg }_{G_{n_{i}}}v_{(\ell)}$ denotes the $\ell$th highest degree in $G_{n_{i}}$. Then, for an unseen graph $G_{n_{j}}$ with $n_{j}$ nodes we estimate the $\ell$th highest degree using
@@
as a means of comparison. Figures 12 and 12 show the results of the four experiments using equations 1 and (151) respectively.
@@ F.1.2 Estimating $U$ with finite partitions
i.e., the ratio of the $j$th highest degree to the sum of all degrees. Figures 14 and 14 show the results of the 4 experiments using equations (4) and (152) respectively.
@@ F.1.3 Estimating $U$ with infinite partitions
For each experiment we use a set of graphs and using equation (5) we estimate $\hat{k}_{i}$, which are 30, 23, 30 and 4 for the four experiments. Figures 16 and 16 show the estimated and actual $p_{j}$ values using equations (4) and (152) respectively.
@@
F.2 Physics citations dataset
We use the physics citations Hep-PH dataset (Leskovec et al. 2005) where nodes are papers and if paper $i$ cites paper $j$, then node $i$ is connected to node $j$. We analyze the growing citation network for months 75, 80, 85, 90, 95 and 100. In each month new papers are published and as such new nodes and edges are added to the the network. For each month $\ell$, we consider $G_{n_{\ell}}$ as the training graph $G_{n_{\ell+j}}$ for $j\in\{5,6,7,8,9,10\}$ as the test graph. We call $j$ the prediction step. Using equation (1) we predict the top 10 degrees of each test graph. Figures 18 and 18 show the results computed using equations 1 and (151), where equation (151) is used as a baseline. Figure 22 shows the absolute errors $|\text{deg }_{G_{n_{j}}}v_{(\ell)}-\text{deg }_{G_{n_{j}}}\hat{v}_{(\ell)}|$ grouped by the prediction step for both the proposed method and the baseline.
Figure 17: The top-10 predicted and actual degrees for the Hep-PH dataset using training graphs $G_{n_{\ell}}$ for $\ell\in\{75,80,85,90,95,100\}$ and test graphs $G_{n_{\ell+j}}$ for $j\in\{5,6,7,8,9,10\}$. The top-10 degrees of each test graph are predicted using equation (1). Different sizes are used to depict the prediction step $j$ and colors are used to depict the degree ranks. The line $y=x$ is drawn to illustrate how close the predicted values are to the actual values.
Figure 18: The top-10 predicted and actual degrees for the Hep-PH dataset using training graphs $G_{n_{\ell}}$ for $\ell\in\{75,80,85,90,95,100\}$ and test graphs $G_{n_{\ell+j}}$ for $j\in\{5,6,7,8,9,10\}$. The top-10 degrees of each test graph are predicted using equation (151). Different sizes are used to depict the prediction step $j$ and colors are used to depict the degree ranks. The line $y=x$ is drawn to illustrate how close the predicted values are to the actual values.
F.3 Social network dataset
Using Facebook (FB) links dataset (Viswanath et al. 2009) we consider training graphs $G_{n_{\ell}}$ on days $\ell\in\{500,520,540,560,580,600\}$. We consider prediction steps $j\in\{50,60,70,80,90,100\}$ and predict the top-10 degrees in test graphs $G_{n_{\ell+j}}$ for $j$ in the set above. Figures 20 and 20 show predicted degrees using equations (1) and (151) respectively. Figure 22 shows the absolute errors $|\text{deg }_{G_{n_{j}}}v_{(\ell)}-\text{deg }_{G_{n_{j}}}\hat{v}_{(\ell)}|$ grouped by the prediction step for both the proposed method and the baseline.
Figure 19: The top-10 predicted and actual degrees for the FB dataset using training graphs $G_{n_{\ell}}$ for $\ell\in\{500,520,540,560,580,600\}$ and test graphs $G_{n_{\ell+j}}$ for $j\in\{50,60,70,80,90,100\}$. The top-10 degrees of each test graph are predicted using equation (1). Different sizes are used to depict the prediction step $j$ and colors are used to depict the degree ranks. The line $y=x$ is drawn to illustrate how close the predicted values are to the actual values.
Figure 20: The top-10 predicted and actual degrees for the FB dataset using training graphs $G_{n_{\ell}}$ for $\ell\in\{500,520,540,560,580,600\}$ and test graphs $G_{n_{\ell+j}}$ for $j\in\{50,60,70,80,90,100\}$. The top-10 degrees of each test graph are predicted using equation (151). Different sizes are used to depict the prediction step $j$ and colors are used to depict the degree ranks. The line $y=x$ is drawn to illustrate how close the predicted values are to the actual values.
Figure 21: Absolute errors $|\text{deg }_{G_{n_{j}}}v_{(\ell)}-\text{deg }_{G_{n_{j}}}\hat{v}_{(\ell)}|$ for Hep-PH dataset grouped by the prediction step for proposed method and the baseline.
Figure 22: Absolute errors $|\text{deg }_{G_{n_{j}}}v_{(\ell)}-\text{deg }_{G_{n_{j}}}\hat{v}_{(\ell)}|$ for FB dataset grouped by the prediction step for proposed method and the baseline.

### v2

@@ B.3 Generating dense and sparse graph sequences with a $(U,W)$ mixture
Figure 13: The dense part $G_{d_{i}}$, sparse part $G_{s_{i}}$ and the mixture graph $G_{n_{i}}$ Both dense part $G_{d_{i}}$ and sparse part $G_{s_{i}}$ have 100 nodes each.
Figure 14: Degree distribution when $\frac{n_{s_{i}}}{n_{d_{i}}}=1$
@@
When $\frac{n_{s_{i}}}{n_{d_{i}}}\to c\in\mathbb{R}$ the graph sequence $\{G_{n_{i}}\}_{i}$ is dense (Lemma 3.4), i.e., the edge density of $G_{n_{i}}$ does not go to zero. Figure 14 shows an example of the dense part $G_{d_{i}}$, the sparse part $G_{s_{i}}$ and the mixture graph $G_{n_{i}}$ for $n_{d_{i}}=n_{s_{i}}=100$. Figure 14 shows the degree distributions for different values of $n_{d_{i}}$ and $n_{s_{i}}$ when the ratio $\frac{n_{s_{i}}}{n_{d_{i}}}=1$. When the ratio $\frac{n_{s_{i}}}{n_{d_{i}}}$ converges to $c\leq 1$, depending on $W$ and $U$, the largest degree may be contributed by the dense part. In Figure 14 the largest degree is actually produced by $U$, however, it is so close to the degrees generated by $W$ that we cannot distinguish the effect of $U$ in this example.
Figure 18 shows the dense part $G_{d_{i}}$ and the sparse part $G_{s_{i}}$ when $n_{d_{i}}=100$ and $n_{s_{i}}=300$. Figure 18 shows the degree distribution when $\frac{n_{s_{i}}}{n_{d_{i}}}=3$. When the ratio $\frac{n_{s_{i}}}{n_{d_{i}}}$ converges to $c>1$, for large enough $i$ the sparse part contributes to highest-degree nodes, which are anomalous. In Figure 18 two anomalous nodes are contributed by $U$ for each pair of $n_{s_{i}}$ and $n_{d_{i}}$.
Figure 15: The dense part $G_{d_{i}}$, sparse part $G_{s_{i}}$ and the mixture graph $G_{n_{i}}$ with $n_{d_{i}}=100$ and $n_{s_{i}}=300$.
@@
Figure 17: The dense part $G_{d_{i}}$, sparse part $G_{s_{i}}$ and the mixture graph $G_{n_{i}}$ with $n_{d_{i}}=300$ and $n_{s_{i}}=1162$.
@@
If $\{G_{n_{i}}\}_{i}$ is sparse, then $n_{s_{i}}$ grows faster than $n_{d_{i}}$ making $\frac{n_{s_{i}}}{n_{d_{i}}}$ go to infinity (Lemma 3.4). We used $n_{s_{i}}=\lceil\sqrt{5i}\times n_{d_{i}}\rceil$ for this experiment and Figure 18 shows the degree distributions. Two anomalous nodes contributed by $U$ can be seen in 3 of the 4 subplots. Figure 18 shows the dense part $G_{d_{i}}$, the sparse part $G_{s_{i}}$ and the mixture graph $G_{n_{i}}$ when $n_{d_{i}}=300$ and $n_{s_{i}}=1162$. We see that the effect of the stars is higher in this mixture.
@@ B.4 Expectations
B.5 Edge density of $(U,W)$-mixture graphs
@@
Let $\{G_{n_{i}}\}_{i}$ be a sequence of sparse $(U,W)$-mixture graphs (Definition 3.2) with dense and sparse parts of $G_{n_{i}}$ being $G_{d_{i}}$ and $G_{s_{i}}$ respectively. Let $\bm{p}=(p_{1},p_{2},\ldots)$ be the mass-partition (Definition 3.1) associated with $U$ and suppose $\bm{p}$ has $k$ non-zero elements. Recall $G_{s_{i}}$ is a union of disjoint stars and isolated edges. Let $\tilde{q}_{j,i}$ be the degree of the star in $G_{s_{i}}$ corresponding to $p_{j}\neq 0$. Then, if we consider $G_{n_{i}}$ to be the disjoint union of $G_{d_{i}}$ and $G_{s_{i}}$ the top $k$ degrees denoted by $\text{deg }v_{(1\ldots k)}$ satisfies
@@
giving the result.
B.6 When $\sum_{i}p_{i}<1$
@@ D.2 Estimating $k$
[…]
@@
Let $\{G_{n_{i}}\}_{i}$ be a sequence of sparse $(U,W)$-mixture graphs (Definition 3.2) with dense and sparse parts $G_{d_{i}}$ and $G_{s_{i}}$ respectively. Let $\text{max\_deg}_{\widetilde{G}_{d_{i}}}$ denote the maximum degree in $G_{n_{i}}$ restricted to nodes from $G_{d_{i}}$. That is, $\text{max\_deg}_{\widetilde{G}_{d_{i}}}$ belongs to a node from the dense part $G_{d_{i}}$. Then
@@
for $\ell\in\{1,\ldots,k-1\}.$ Let $\text{max\_deg}_{\widetilde{G}_{d_{i}}}$ denote the maximum degree in $G_{n_{i}}$ restricted to nodes from $G_{d_{i}}$. That is, $\text{max\_deg}_{\widetilde{G}_{d_{i}}}$ belongs to a node from the dense part $G_{d_{i}}$. For $W=1$ we have $\text{max\_deg}_{\widetilde{G}_{d_{i}}}=n_{d_{i}}-1$ and
@@
But for all other $W\neq 1$, $\text{max\_deg}_{\widetilde{G}_{d_{i}}}<<n_{d_{i}}$. For large enough $i$, (k+1)(k+1)st highest degree
@@
[…]
@@ Appendix E Estimating $U$ for infinite partitions
We turn out attention to the term $\hat{k}_{i}$ . We note that $\hat{k}_{i}\in O(m_{s_{i}}/n_{d_{i}})$. The sequence $\{\hat{k}_{i}\}_{i}$ increases at a slower rate compared to $m_{s_{i}}/n_{d_{i}}$ depending on the rate $\{p_{j}\}_{j}$ decreases. For $\sum_{j}p_{j}$ to add to 1, $p_{j}$ cannot decrease like $c/j$ for come constant $c$, because $\sum_{j}c/j$ diverges to infinity. Hence $p_{j}\approx c/j^{1+\alpha}$ is an option. The estimate $\hat{k}_{i}$ satisfies
@@
$\lim_{i\to\infty}\frac{1}{m_{s_{i}}}\text{Var}(Y_{i})=\sum_{\ell=1}^{\hat{k}_{i}}p_{\ell}\left(1-\sum_{\ell=1}^{\hat{k}_{i}}p_{\ell}\right)\,.\\$
@@
$\frac{1}{(j+1)^{1+\alpha}}$ $<\int_{j}^{j+1}\frac{1}{x^{1+\alpha}}\,dx\,$, (148)
$\sum_{j=\hat{k}_{i}}^{\infty}p_{j}<\sum_{j=\hat{k}_{i}}^{\infty}\frac{1}{(j+1)^{1+\alpha}}$ $\leq\int_{\hat{k}_{i}}^{\infty}\frac{1}{x^{1+\alpha}}\,dx\,$, (149)
@@ F.1.1 Degree prediction
[…]
@@
as a means of comparison. Figures 22 and 22 show the results of the four experiments using equations 1 and (151) respectively.
@@ F.1.2 Estimating $U$ with finite partitions
i.e., the ratio of the $j$th highest degree to the sum of all degrees. Figures 24 and 24 show the results of the 4 experiments using equations (4) and (152) respectively.
@@ F.1.3 Estimating $U$ with infinite partitions
For each experiment we use a set of graphs and using equation (5) we estimate $\hat{k}_{i}$, which are 30, 23, 30 and 4 for the four experiments. Figures 26 and 26 show the estimated and actual $p_{j}$ values using equations (4) and (152) respectively.
@@
F.2 Real datasets
We use Facebook (FB) links (Viswanath et al. 2009), Hep-PH Physics citations (Leskovec et al. 2005, Gehrke et al. 2003), MOOC interactions (Kumar et al. 2019), SMS (Wu et al. 2010), UCI Messages (Panzarasa et al. 2009) and Yahoo messages (Rossi & Ahmed 2015) datasets. Each dataset is given as an edgelist with timestamped edges. The edges accumulate at different speeds in each dataset. In Yahoo messages dataset the edges accumulate quite fast and as such we use a 2-hourly time window to construct graphs. In contrast, edges accumulate slowly in Hep-PH citation dataset and we use monthly graphs. For the other datasets we use daily graphs. In each example we consider growing networks, i.e., new edges and nodes are added to the existing graph as time passes.
Each dataset comprises a sequence of growing graphs $\{G_{n_{i}}\}_{i}$ and we consider $G_{n_{i}}$ for $i\in\{20,\ldots,24\}$ as training graphs. For each $G_{n_{i}}$ we select the test graph $G_{n_{j}}$ such that $j=\min_{k}\{|G_{n_{k}}|:|G_{n_{k}}|-|G_{n_{i}}|\geq 500\}$, i.e., $G_{n_{j}}$ is the first graph that has more than 500 nodes compared to $G_{n_{i}}$. This is to ensure that the test graph is different from the training graph. Using the training graphs we predict the top-10 degrees of the test graphs.
For Bollobás et al. (2001) we used equation (151) to predict degrees. Caron & Fox (2017) have made their code available at https://www.stats.ox.ac.uk/˜caron/code/bnpgraph/. We used their code to train their graph generation model on the training graph $G_{n_{i}}$ and using the trained model generated a graph having $n_{k}$ nodes, which we compared with $G_{n_{k}}$. Kronecker graphs have 2 settings: deterministic and stochastic, and their code is available at https://github.com/snap-stanford/snap/tree/master/examples/krongen. We generated Kronecker graphs with (1) the default setting (stochastic $2\times 2$ initiator matrix), (2) stochastic setting with a $3\times 3$ initiator matrix and (3) the deterministic setting with a $3\times 3$ initiator matrix. For the stochastic settings, we first fitted a Kronecker model on the training graph $G_{n_{i}}$ and then using the fitted parameters generated the graph with $n_{k}$ nodes, which we compared with the test graph $G_{n_{k}}$. Leskovec et al. (2010) explored the initiator matrix resulting from a 3-star ($K_{1,3}$) with self loops added at each node. The adjacency matrix of this graph is given by
$A=\begin{pmatrix}1&1&1\\ 1&1&0\\ 1&0&1\end{pmatrix}\,$. (153)
We explored the deterministic case with this initiator matrix. Using the predicted top-$10$ degrees from each method, we computed the average MAPE and the standard deviation, which is given in Table 2. We compared the performance of the top 2 methods on each dataset using the Student’s t-test with $\alpha=0.1$. For all datasets except UCI Messages we found that $(U,W)$-mixture was significantly better than Bollobás et al. (2001), which was the ranked the second. For UCI, while $(U,W)$-mixtures performed better on average, it was not significantly better. Figures 32 to 32 show the actual and predicted degrees of the test graphs with Actual = Predicted line drawn. In some instances, the $y=x$ line is out of range.
Figure 27: Facebook links dataset
Figure 28: HEP-PH dataset
Figure 29: MOOC dataset
Figure 30: UCI Messages dataset
Figure 31: SMS dataset
Figure 32: Yahoo dataset
F.2.1 HEP-TH dataset comparison as in Leskovec et al. (2005)
Leskovec et al. (2010) have used the initiator matrix $A$ given in equation (153) to generate graphs similar to the HEP-TH dataset using the deterministic setting. For the stochastic setting, they have replaced 1s with $\alpha=0.41$ and 0s with $\beta=0.11$. With these two initiators we ran their algorithm and verified we could produce the degree distribution in their paper. In their paper they say that Kronecker graph results qualitatively match the original results. We found this to be true in our simulation as well. Table 6 gives the actual and predicted top-10 degrees using $(U,W)$-mixture graphs, Kronecker deterministic and stochastic versions and the absolute percentage error $|\hat{y}-y|\times 100/y$.
Table 6: The top-10 degree comparison results on HEP-TH dataset where APE denotes the absolute percentage error $|\hat{y}-y|\times 100/y$.
Actual Predicted $(U,W)$ degree Predicted Kronecker Deterministic. Predicted Kronecker Stochastic APE $(U,W)$ APE Kronecker Det. APE Kronecker Stoch.
2468 2249 32768 74 8.89 1227.71 97.00
1797 1633 16384 51 9.14 811.74 97.16
1653 1503 16384 47 9.09 891.17 97.16
1369 1281 16384 45 6.44 1096.79 96.71
1308 1253 16384 44 4.19 1152.60 96.64
1219 1147 16384 44 5.92 1244.05 96.39
1218 1124 16384 43 7.69 1245.16 96.47
1165 1101 16384 42 5.51 1306.35 96.39
1124 1019 16384 42 9.35 1357.65 96.26
1038 996 16384 42 4.01 1478.42 95.95